Given that the Fermi energy of gold is 5.54 eV, the number density of electrons is? ( use me=9.11×10−31 kg; h=6.626×10−34Js; 1eV=1.6×10−19J )
1
Expert's answer
2021-09-23T08:31:26-0400
"n=\frac{2m\pi}{h^2} E_F=\\\space\\
=\frac{2(9.11·10^{-31})·3.14}{(6.626·10^{−34})^2}·5.54·1.6·10^{-19}=\\\space\\
=2.31·10^{19}\text{ m}^{-3}."